In a study of pain relievers,

100

people were given product A, and all but

23

experienced relief. In the same study,

50

people were given product B, and all but

8

experienced relief.

Fill in the blanks of the statement below to make the statement the most reasonable possible.

Product ?AB performed better in the study because only

%

failed to get relief with this product, whereas

%

failed to get relief with Product ?AB
Answer #1

product A = 23/100 = 0.23

ProductB = 8/50 = 0.16

Product A performed better in the study because only
77% failed to get relief with this product, whereas
84% failed to get relief with Product B